A transfer shaft with two spur gears shown below: Input Gear 1 Output Gear Torque Bearing 1 Gear 1 Gear 2 Bearing 2 B OTO 4Slopes 0.0005-0.0012 rad 0.0008-0.0012 rad Tapered roller Cylindrical roller Deep-groove ball Spherical ball Self-align ball1. Diameter of the shaft required by consideration of the slope at the bearings? 2. Diameter of the shaft required by conside
Answer #1

The free body diagram for the system on the horizontal and vertical plane is prepared. Then, Shear Force and Bending Moment diagram is prepared for horizontal and vertical planes. The maximum combined stress due to torque and moment is computed at two points and then mean and amplitude stress are computed in terms of d. The Ka, Kb, Kc, Kd are computed using the tables and the data provided (Refer Shigley). Once Ka, Kb, Kc and Kd are known, the diameter is evaluated using the Goodman Criterion for fatigue failure with a factor of safety of 2 (as given in the problem). Using the given torsional wind up, the diameter is calculated by evaluating torsional rigidity as the Young's Modulus and poisson's ratio are given and the torque acting is known.

Note: The problem is lengthy and whatever was possible to solve in the given time is done. It's difficult to solve all the parts in two hours.

T= 2500 lbf-in T= 2500 llaf-in M = 1200+ 4 = 1204 lbf-in Sport (939) = (17954.500) M = 7200 + 16 = 7216 lbf-in 16 (M+ M² + 2)